# Weak continuous measurements require more work than strong ones

### v2

@@ Abstract
We analyze a dynamical model for a quantum measurement process, able to capture nonideal (weak or inefficient) measurements. In this model, the irreversibility of the measurement dynamics is due to action of a reservoir at equilibrium, which cause decoherence between the states associated with different measurement results. We analyze the performance of measurement process generated by the model, by introducing figures of merits to quantify the strength of the measurement and its efficiency. We also derive and analyze a lower bound on the measurement work cost that we can relate to the measurement quality. We take as an illustration a qubit measurement owing to its coupling to a harmonic oscillator. We investigate the long sequences of extremely short and weak measurements (a.k.a continuous measurements), to find under which conditions they converge to an ndeal measurement and analyze their work cost. Surprisingly, we find that a sequence converging towards a projective measurement has a much larger work cost than a equivalent strong measurement obtained from a single intense interaction with the apparatus. We extend this result to a large class of models owing to scaling arguments. Our analysis offer new insights into the trade-offs between measurement strength, energy consumption, and information extraction in quantum measurement protocols.
@@ I Introduction
Quantum measurements are among the most intriguing processes in quantum mechanics and have been the subject of extensive discussion for decades [1, 2, 3]. Quantum theory allows for various types of measurements, ranging from the simplest—projective measurements—to more sophisticated techniques such as POVMs (Positive Operator-Valued Measures), quantum non-demolition measurements, and homodyne detection [4, 5, Jordan24]. In practical experiments, measurement protocols are non-ideal due to factors such as detector inefficiency, thermal noise, imperfections in state preparation and finite interaction time with the measured system. These sources of non-ideality can significantly impact the measurement process, and their effects can be modeled and quantified [6, 7, 5].
@@
All quantum measurement processes share the fundamental requirement to generate a family of macroscopically distinguishable states of the apparatus, that can be read owing to conventional (non quantum) apparatuses, yielding a measurement result. The latter is a classical stochastic variable whose value can be stored into a classical memory device. The transition towards a classical behavior of the measured system and the degrees of freedom of the apparatus encoding the measurement result inherently involves a non-unitary process [8, 9]. It can therefore be derived from a purely unitary evolution of the system and the apparatus only if the latter contains many degrees of freedom, besides the ones accessible to the observer, over which a trace can be taken. This requirement is closely related to the theory of quantum Darwinism, which posits that the environment facilitates the emergence of classicality by encoding redundant copies of information (i.e. of the measurement result), thereby making it accessible to multiple observers, and hence an objective fact [10].
@@
Since quantum measurements cannot be modeled by a purely unitary process, they constitute a non-trivial thermodynamic transformation, involving exchanges of entropy and energy. These thermodynamic resources have been extensively explored, notably in the context of powering quantum engines and cooling mechanisms [11, 12, 13, 14, 15]. Despite the central role of quantum measurements in virtually all quantum technologies and applications [16, 17, 18], relatively few studies have addressed the work cost associated with performing quantum measurements [19, 20, 21, 22, 23, 24, 25]. A key result found in the literature is that ideal projective measurements, in principle, require infinite resources, making it impossible to achieve them exactly. Instead, they can only be (arbitrarily well) approximated [26].
@@
One might ask what the minimum amount of work required for a quantum measurement is, and what the optimal protocol would be to achieve this minimum. In a previous study, some of us derived a lower bound on the work cost required to perform a quantum measurement in terms of figure of merit quantifying the information acquired by the measurement [25]. To obtain such a result, the measuring apparatus was modeled as a pointer continuously coupled to a reservoir able to bring it back to thermal equilibrium. The obtained lower bound was then based on the expression of the Second Law of thermodynamics during the measurement process. Deviations from the minimum work cost was associated with the entropy produced when operating and resetting the measurement device. Surprisingly, a thermodynamically reversible measurement protocol, achieving asymptotically the minimum cost, was identified and analyzed. In this work, we investigate two new questions: First, is it optimal to achieve classicality owing to a thermalizing bath (in which energy can be wasted under the form of heat dissipation), or is a pure dephasing process (comprising no heat transfer) preferable? We find that a pure dephasing bath actually leads to a less favorable lower bound on the work cost. Second, how is the energy cost scaling with the strength of the measurement (the amount of extracted information): In other words, is it more energetically favorable to perform a sequence of weak measurements converging towards a projective measurement [27], or a single nearly-projective measurement? Our analysis shows that the converging sequence of weak measurements requires much more energy in general.
@@
The paper is organized as follows: First, we introduce in Sec. II the quantum measurement protocol considered throughout the paper. Next, we examine in Sec. III the energetics of the dephasing process and derive the lower bound on the measurement work cost. Sec. IV focus on a specific model involving a qubit and a harmonic oscillator (cavity). In Sec. (V), we introduce and discuss measures of non-ideality, specifically strength and efficiency of the measurement. We use them in Sec. VI to analyze the work cost and the performance of a single qubit measurement, and in Sec. VII to compare long sequences of weak continuous measurements with single strong measurement.
@@
Figure 1: Measurement protocol: (0) ancilla state preparation, (1) ancilla-system interaction, (2) ancilla dephasing, and (3) ancilla readout, (4) ancilla reset and (5) memory reset. The color change in both the system and ancilla indicates a change of state. Blurred disks represent quantum states, while filled disks denote classical states.
@@ II Dynamical model of quantum measurement
In this article we consider the following measurement protocol involving a system S in state $\hat{\rho}_{S}$ (the measured system) and an ancilla A, initialized at time $t_{0}$ in state $\hat{\rho}_{A}(t_{0})$ (Step (0) in Fig. 1). The ancilla A models both the part of the apparatus which first interacts with S, as well as the degree of freedom in which the measurement result will be read by the observer. In contrast, the influence of other inaccessible degrees of freedom of the apparatus will be included via the non-unitary dynamics they induce on A. The combined system-ancilla state $\hat{\rho}_{SA}\left(t_{0}\right)=\hat{\rho}_{S}(t_{0})\otimes\hat{\rho}_{A}(t_{0})$ is initially separable, meaning no information about S is a priori contained in A. Next, the system and ancilla unitarily interact, transferring information from the system to the ancilla (step (1) in Fig. 1). The interaction lasts for $\Delta t=t_{1}-t_{0}$ and is governed by the unitary operator $\hat{U}(\Delta t)$, which possibly also includes control operations affecting only the ancilla. The full system-ancilla state after the interaction is
$\hat{\rho}_{SA}\left(t_{1}\right)$ $=$ $\hat{U}\left(\Delta t\right)\left(\hat{\rho}_{S}\otimes\hat{\rho}_{A}\left(t_{0}\right)\right)\hat{U}^{\dagger}\left(\Delta t\right).$ (1)
@@
The amount of correlations that are generated between the system and the ancilla (or equivalently, the amount of information about the system extracted during the interaction) is set by the coupling constant and the interaction time $\Delta t$: in the case of a weak interaction, or short interaction time, only a small amount of information is transferred, and probing the ancilla will only weakly affect the system state. Conversely, a strong/long enough interaction time may fully copy in the ancilla all the information about the system observable involved in the interaction (that is, generate maximal correlations between the ancilla and the system states), leaving the system in a fully mixed reduced state [28].
@@
Following the interaction step, we assume that the ancilla undergoes a strong pure dephasing process (step (2) in Fig. 1). This step models the “objectification process”, during which the information about the system stored in the ancilla is copied multiple times into a macroscopic amount of degrees of freedom, making the measurement result an objective fact [10, 29, 25].
In this step, a basis of stable (pointer) states of the ancilla is selected by the non-unitary environment-induced dynamics, while coherences between those pointer states are completely suppressed. This process can model internal dynamics of a macroscopic system composed of interacting degrees of freedom [30] – the inaccessible degrees of freedom of the apparatus–, or irreversible processes taking place when the signal from the apparatus is extracted and amplified. To analyze the differences with the protocol analyzed in [25], we restrict ourselves in the following to a bath inducing pure dephasing in the energy eigenbasis $\{{\left|{n}\right\rangle}\}$ of the ancilla. Pure-dephasing in another basis can effectively be achieved in engineered environments resulting from driven-dissipative processes (e.g. in a quantum limited amplifier, [31]). We, however, leave the analysis of such a case for another study, and focus here on minimal ingredients enabling to describe a quantum measurement. The fact that pure dephasing starts only after the system-ancilla interaction took place captures experimentally relevant situations when the interaction and dephasing scales are very different, or when the pure dephasing occurs during a later-stage amplification process.
@@
Furthermore, we require that this objectification process encompasses any evolution with nontrivial energetic and thermodynamic role involved in the measurement process, by imposing that at the end of step (2), the ancilla is fully dephased, i.e. in a classical mixture of the energy eigenstates:
$\hat{\rho}_{SA}\left(t_{f}\right)$ $=$ $\sum_{n}{\left|{n}\right\rangle}{\left\langle{n}\right|}\hat{U}\left(\Delta t\right)\left(\hat{\rho}_{S}\otimes\hat{\rho}_{A}\left(t_{1}\right)\right)\hat{U}^{\dagger}{\left|{n}\right\rangle}{\left\langle{n}\right|}$
@@
$p_{n}$ $=$ $\text{Tr}\left[\left(\hat{I}\otimes\hat{\Pi}_{n}\right)\hat{U}\left(\Delta t\right)\left(\hat{\rho}_{S}\otimes\hat{\rho}_{A}\right)\hat{U}^{\dagger}\left(\Delta t\right)\left(\hat{I}\otimes\hat{\Pi}_{n}\right)\right].\;\;\;\;$ (2)
where $\hat{\rho}_{S|n}$ (resp. $p_{n}$) is the conditional system state associated with (the probability) finding the ancilla in state ${\left|{n}\right\rangle}$.
Finally, the ancilla is read out in the basis set by the dephasing process, here the energetic eigenbasis ((3) in Fig. 1). We assume that the dephasing process singles out a set of macroscopically distinguishable states of the ancilla, which can be read via a classical measurement (either directly in the ancilla if it is a macroscopic degree of freedom, like a large-amplitude field mode, or indirectly via its impact in the environment, for instance via the light it emits). As we required the dephasing to be complete during the objectification step, this process only corresponds to information acquisition, without exchange of energy. Its only impact on the quantum state describing the system and ancilla is the selection of an element of the classical mixture, just as a measurement performed on a probabilistic ensemble of classical systems. While in principle all the ${\left|{n}\right\rangle}$ can be perfectly discriminated (being orthogonal quantum states), it is more reasonable to assume that macroscopically distinguishable observables will correspond to coarse-grained subsets of those states. A convenient way to introduce such subsets is via a set of orthogonal projectors $\hat{\Pi}_{r}=\sum_{n}p(r|n)\hat{\Pi}_{n}$, where $\hat{\Pi}_{n}={\left|{n}\right\rangle}{\left\langle{n}\right|}$ and where the index $r$, the measurement outcome, labels the distinguishable coarse-grained subspaces. We furthermore require $\sum_{r}\hat{\Pi}_{r}=\mathds{1}$ and $\hat{\Pi}_{r}\hat{\Pi}_{r^{\prime}}=\delta_{r,r^{\prime}}\hat{\Pi}_{r}$ such that $p(r|n)$ is a piecewise-constant function with values in $\{0,1\}$. Injecting the latter relation in equation (II) leads to:
@@
and $p_{r}=\sum_{n}p_{n}p(r|n)$ the probability to obtain the outcome $r$.
After reading the ancilla’s state and storing the information in a classical memory, the measurement process is practically over. However, to make sure all thermodynamic costs are included in our protocol, we also include the reset of the ancilla and of the memory (after its information content has been potentially used for some task) to their initial states ((4 & 5) in Fig. 1.
This measurement protocol can model a wide range of quantum measurements, from strong to weak and from ideal to noisy. Like the protocol analyzed in [25], it does not involve any black-box quantum measurement on the ancilla which would make the thermodynamic analysis incomplete.

### v3

@@ Abstract
Understanding the energy cost of quantum measurement process and its connection to the measurement performance faces the challenge of modeling the objectification process. The latter, turns the measurement result into an objective fact, available to independent observers, and is responsible for the measurement irreversibility. To address this issue, we propose and analyze a dynamical model of quantum measurement, able to capture nonideal (weak and inefficient) measurements. In this model, the objectification is induced by a contact with a macroscopic reservoir at equilibrium which is responsible for the redundant broadcast of the measurement outcome (producing a Spectrum Broadcast Structure (SBS) state) while inducing decoherence in the pointer basis, in the line of the theory of quantum Darwinism. We analyze the performance of the obtained measurement process by introducing figures of merit to quantify the strength of the measurement and its efficiency. We also derive and a lower bound on the measurement work cost that we can relate to the measurement quality. We take as an illustration the readout of a qubit via its coupling to a harmonic oscillator. We investigate the long sequences of extremely short and weak measurements (a.k.a continuous measurements), to find under which conditions they converge to an ideal (projective) measurement and analyze their work cost. Surprisingly, we find that a sequence converging to projective measurement has a much larger work cost than an equivalent strong measurement obtained from a single intense interaction with the apparatus. We extend this result to a large class of models owing to scaling arguments. Our analysis offers new insights into the trade-offs between measurement strength, energy consumption, and information extraction in quantum measurement protocols.
@@ I Introduction
Quantum measurements are among the most intriguing processes in quantum mechanics, and their modeling has been the subject of extensive discussion for decades [braginsky1995quantum, Allahverdyan13, brukner2017quantum, griffiths2017quantum]. Quantum theory allows for various types of measurements ranging from the ideal projective measurements - introduced by the measurement postulate of quantum mechanics - to more sophisticated processes, such as POVMs (Positive Operator-Valued Measures), quantum non-demolition measurements, and homodyne detection [busch2016quantum, wiseman2009quantum, Jordan242]. In practical experiments, measurement protocols are non-ideal due to factors such as detector inefficiency, thermal noise, imperfections in state preparation and finite interaction time with the measured system. These sources of non-ideality can significantly impact the measurement process, and their effects can be modeled and quantified [auffeves2019generic, warszawski2002quantum, wiseman2009quantum].
@@
All quantum measurement processes share the fundamental requirement to generate a family of macroscopically distinguishable states of the apparatus, that can be read owing to conventional (non quantum) apparatuses, yielding a measurement result (or outcome). The latter is a classical stochastic variable whose value can be stored into a classical memory device. The transition towards a classical behavior of the degrees of freedom of the apparatus encoding the measurement result inherently involves a non-unitary process [neumann1955mathematical, joos2013decoherence]. It can therefore be derived from a purely unitary evolution of the system and the apparatus only if the latter contains many degrees of freedom, in addition to the ones accessible to the observer, over which a trace can be taken. This requirement is closely related to the theory of quantum Darwinism, which posits that the environment facilitates the emergence of classicality by encoding redundant copies of information (i.e. of the measurement result), thereby making it accessible to multiple observers and hence an objective fact [zurek2009quantum].
@@
Since quantum measurements cannot be modeled by a purely unitary process of the system and a pointer, they constitute a non-trivial thermodynamic transformation, involving exchanges of entropy and energy. These thermodynamic resources have been extensively explored, notably in the context of powering quantum engines and cooling mechanisms [parrondo2015thermodynamics, elouard2018efficient, buffoni2019quantum, Jordan20, Fadler23]. Despite the central role of quantum measurements in virtually all quantum technologies and applications [fellous2023optimizing, pirandola2020advances, degen2017quantum], relatively few studies have addressed the work cost associated with performing quantum measurements [Jacobs09, Sagawa2009Jun, deffner2016quantum, debarba2019work, Mohammady2023Jan, Piccione2023Nov, latune2024thermodynamically]. A key result found in the literature is that ideal projective measurements, in principle, require infinite resources, making it impossible to achieve them exactly. Instead, they can only be (arbitrarily well) approximated [guryanova2020ideal].
@@
One might ask what the minimum amount of work required for a quantum measurement is and what the optimal protocol would be to achieve this minimum. In a previous study, some of us derived a lower bound on the work cost required to perform a quantum measurement in terms of figures of merit quantifying the information acquired by the measurement [latune2024thermodynamically]. To obtain such a result, the measuring apparatus was modeled as a pointer continuously coupled to a reservoir able to bring it back to thermal equilibrium. The obtained lower bound was then based on the expression of the Second Law of thermodynamics during the measurement process. Deviations from the minimum work cost were associated with the entropy produced when operating and resetting the measurement device. Surprisingly, a thermodynamically reversible measurement protocol, achieving asymptotically the minimum cost, was identified and analyzed. However, this setup could not capture the paradigmatic case of a weak (incomplete) but efficient (purity-preserving) measurement [jacobs2006straightforward].
In this work, we investigate two new questions: First, is it optimal to achieve classicality owing to a thermalizing bath (in which energy can be wasted under the form of heat dissipation), or is a pure dephasing process (comprising no heat transfer) preferable? Surprisingly, we find that a pure dephasing bath actually leads to a less favorable lower bound on the work cost. Second, how is the energy cost scaling with the strength of the measurement (the amount of extracted information): In other words, is it more energetically favorable to perform a sequence of weak measurements converging towards a projective measurement [Jacobs2006Sep], or a single nearly-projective measurement? Our analysis shows that, despite the work cost of the measurement vanishing with the measurement strength, the converging sequence of weak measurements requires much more energy in general, revealing a previously unrecognized fundamental trade-off, which constitutes the central result of this work.
@@
The paper is organized as follows: First, we introduce in Sec. II the quantum measurement protocol considered throughout the paper. Next, we examine in Sec. III the energetics of the dephasing process and derive the lower bound on the measurement work cost. Sec. IV focuses on a specific model involving a qubit and a harmonic oscillator (cavity). In Sec. (V), we introduce and discuss measures of non-ideality, specifically strength and efficiency of the measurement. We use them in Sec. VI to analyze the work cost and the performance of a single qubit measurement, and in Sec. VII to compare long sequences of weak continuous measurements with single strong measurement.
@@
Figure 1: Measurement protocol: (0) ancilla state preparation, (1) ancilla-system interaction, (2) ancilla dephasing, (3) ancilla readout, (4) ancilla reset and (5) memory reset. The color change in both the system and ancilla indicates a change of state. Blurred disks represent quantum states, while filled disks denote classical states.
@@ II Dynamical model of quantum measurement
In this article, we consider the following measurement protocol involving a system S in state $\hat{\rho}_{S}$ (the measured system) and an ancilla A, initialized at time $t_{0}$ in state $\hat{\rho}_{A}(t_{0})$ (Step (0) in Fig. 1). The ancilla A models both the part of the apparatus that first interacts with S, as well as the degrees of freedom in which the measurement result will be read by the observer. In contrast, the influence of other inaccessible degrees of freedom of the apparatus will be included via the non-unitary dynamics they induce on A. The combined system-ancilla state $\hat{\rho}_{SA}\left(t_{0}\right)=\hat{\rho}_{S}(t_{0})\otimes\hat{\rho}_{A}(t_{0})$ is initially separable, meaning no information about S is a priori contained in A. Next, the system and the ancilla unitarily interact, transferring information from the system to the ancilla (step (1) in Fig. 1). The interaction lasts for $\Delta t=t_{1}-t_{0}$ and is governed by the unitary operator $\hat{U}(\Delta t)$, which possibly also includes control operations affecting only the ancilla. The full system-ancilla state after the interaction is
$\hat{\rho}_{SA}\left(t_{1}\right)$ $=$ $\hat{U}\left(\Delta t\right)\left(\hat{\rho}_{S}{\color[rgb]{0,0,0}{\left(t_{0}\right)}}\otimes\hat{\rho}_{A}\left(t_{0}\right)\right)\hat{U}^{\dagger}\left(\Delta t\right).$ (1)
@@
The amount of correlations that are generated between the system and the ancilla (or equivalently, the amount of information about the system extracted during the interaction) is set by the coupling constant and the interaction time $\Delta t$: in the case of a weak interaction, or short interaction time, only a small amount of information is transferred, and probing the ancilla will only weakly affect the system state. Conversely, a strong/long enough interaction time may fully copy in the ancilla all the information about the system observable involved in the interaction (that is, generate maximal correlations between the ancilla and the system states), leaving the system in a fully mixed reduced state [jacobs2006straightforward].
@@
Following the interaction step, we assume that the ancilla undergoes a strong pure dephasing process (step (2) in Fig. 1). This step models the “objectification process”, during which the information about the system, stored in the ancilla, is redundantly copied multiple times into independent subsets of macroscopic amount of degrees of freedom, thereby, making the measurement result an objective fact [zurek2009quantum, Mohammady2021Dec, latune2024thermodynamically].
Figure 2: The ancilla undergoes dephasing through multiple independent environments. This occurs via distinct channels, each characterized by a dephasing coefficient $\gamma_{i}$. At the end of the process, the joint state of the ancilla and the complete system is in a Spectrum Broadcast Structure (SBS) state.
This process can arise from internal dynamics of a macroscopic system [Engineer2024Mar] - the inaccessible degrees of freedom of the apparatus - or from the irreversible processes taking place when the signal from the apparatus is extracted and amplified. The objectification process selects a basis of stable (pointer) states of the ancilla associated to the different measurement outcomes, while coherences between those pointer states are completely suppressed.
To model this step, we assume that the ancilla interacts simultaneously with distinct, non-interacting subsets of the macroscopic degrees of freedom (or, in the language of open quantum systems, separate environments), with potentially different dephasing interaction coefficients, as illustrated in Fig. 2. Once dephasing has lasted long enough to fully decohere the ancilla state, the joint ancilla–environments state takes the form of a so-called SBS state [horodecki2015quantum], which ensures that the measurement outcome could be retrieved independently from a measurement in each of the separate sub-environments (See Appendix F for derivation and further details). Pure dephasing, that is, decoherence in the absence of energy exchange with the environment, naturally emerges from such microscopic models whenever the ancilla-reservoir coupling commutes with the ancilla’s Hamiltonian. The resulting dynamics then show the damping of coherences in the ancilla’s energy eigenbasis $\{{\left|{n}\right\rangle}\}$, which therefore plays the role of a pointer basis. Pure-dephasing in another basis can be achieved in engineered environments resulting from driven-dissipative processes (e.g. in a quantum limited amplifier [Flurin2014Thesis]). Those processes amount to temporarily modify the effective Hamiltionian of the ancilla (e.g. owing to an external driving), such that the interaction selects another pointer basis.
@@
Furthermore, we require that this objectification process encompasses any evolution with nontrivial energetic and thermodynamic role involved in the measurement process, by imposing that at the end of step (2), the ancilla is fully dephased, i.e. in a classical mixture of the pointer states:
[…]
@@
$p_{n}=\text{Tr}$ $\Big[\Big($ $\hat{I}\otimes\hat{\Pi}_{n}\Big)\hat{U}\left(\Delta t\right)\left(\hat{\rho}_{S}{\color[rgb]{0,0,0}{\left(t_{0}\right)}}\otimes\hat{\rho}_{A}{\color[rgb]{0,0,0}{\left(t_{0}\right)}}\right)\hat{U}^{\dagger}\left(\Delta t\right)\left(\hat{I}\otimes\hat{\Pi}_{n}\right)\Big].\quad\;$
where $\hat{\rho}_{S|n}$ (resp. $p_{n}$) is the conditional system state associated with (the probability) finding the ancilla in the state $\left|{n}\right\rangle$}.
Finally, the ancilla is read out in the basis set by the dephasing process, here the energetic eigenbasis ((3) in Fig. 1). We assume that the dephasing process singles out a set of macroscopically distinguishable states of the ancilla, which can be read via a classical measurement (either directly in the ancilla if it is a macroscopic degree of freedom, like a large-amplitude field mode, or indirectly via its impact in the environment, for instance via the light it emits). As we required the dephasing to be complete during the objectification step, this process corresponds only to information acquisition, without exchange of energy. The only impact of the readout process in the quantum state describing the system and the ancilla is the selection of an element of the classical mixture, just as a measurement performed on a probabilistic ensemble of classical systems. While in principle all $\left|{n}\right\rangle$} can be perfectly discriminated (being orthogonal quantum states), it is more reasonable to assume that macroscopically distinguishable observables will correspond to coarse-grained subsets of those states. A convenient way to introduce such subsets is via a set of orthogonal projectors $\hat{\Pi}_{r}=\sum_{n}p(r|n)\hat{\Pi}_{n}$, where $\hat{\Pi}_{n}={\left|{n}\right\rangle}{\left\langle{n}\right|}$ and where the index $r$, the measurement outcome, labels the distinguishable coarse-grained subspaces. We furthermore require $\sum_{r}\hat{\Pi}_{r}=\mathds{1}$ and $\hat{\Pi}_{r}\hat{\Pi}_{r^{\prime}}=\delta_{r,r^{\prime}}\hat{\Pi}_{r}$ such that $p(r|n)$ is a piecewise-constant function with values in $\{0,1\}$. Injecting the latter relation into Eq. (2) leads to:
@@
and $p_{r}=\sum_{n}p_{n}p(r|n)$ the probability of obtaining the outcome $r$.
After reading the ancilla’s state and storing the information in a classical memory, the measurement process is practically over. However, to make sure all thermodynamic costs are included in our protocol, we also include the reset of the ancilla and the memory (after its information content has been potentially used for some task) to their initial states ((4 & 5) in Fig. 1.
This measurement protocol can model a wide range of quantum measurements, from strong to weak and from ideal to noisy. Like the protocol analyzed in [latune2024thermodynamically], it does not involve any black-box quantum measurement on the ancilla which would make the thermodynamic analysis incomplete.